OpenStudy (anonymous):

what are the x, y slope intercept points for the equation x + 2y = 18

OpenStudy (anonymous):

1. y = –2x + 1 2. y = –2x – 1 3. y = –2x – 3 4. y = –2x + 3

OpenStudy (anonymous):

yes thats correct

OpenStudy (anonymous):

in slope intercept form y=1/2x+9 to find x intercept, set y=0 and solve for x 0=1/2x+9 -9=1/2x x=-18 to find y intercept, set x=0 and solve for y y=9 so your intercept points are (-18,9)

OpenStudy (anonymous):

anyone know the steps to graphing this equation?